Jacksonville Traders manufactures battery operated toy cars that would last 20,000 hours.

Random samples of 14 battery operated toy cars were examined. The standard deviation of the sample was 1750 hours. What is the probability that the toy cards, when examined in a test, will not last more than 19800 hours?
1
Expert's answer
2020-10-05T10:59:42-0400
"20000-1750=18250"


"n=14"


"\sigma =1750"

"p=\frac{19800} {20000}=0.99"


"q=1-p=1-0.99=0.01"